Calculate the elctric field at a distance 'r' from the center of a uniformly charged solid sphere of radius 'R' when r < R. The charge density is given by ρ

ρ. 


1
Expert's answer
2021-03-03T11:25:17-0500

Radius of the sphere =R

The charge density ="\rho"


Charge on the sphere =

"\rho\times V=\frac{4}{3} \pi R^3\times \rho"





For r<R using Gauss law we have,


"E(4 \pi r^2) = \frac{1}{\epsilon 0} \times" "\frac{4}{3} \pi R^3\times\rho"


Simplifying for electric field, we have,


"E=\frac{1}{4\pi \epsilon0} \frac{Q}{R^3} r"